OpenStudy (anonymous):

Please help A second grade class had a spelling bee with 5-letter and 6-letter words. They used 37 words that had a total of 201 letters. Which system of linear equations represents the problem

OpenStudy (anonymous):

Remember that for this type of problem, you can se tup an equation using the number of words. You can also set up an equation for the number of letters. x + y = ? 5x + 6y = ?

OpenStudy (anonymous):

I cant figure out where the numbers go in the equation.

OpenStudy (anonymous):

Okay. Th first equation that I have is the number of words. How many words were used total?

OpenStudy (anonymous):

37

OpenStudy (anonymous):

Ok. x + y = 37 How many letters were used?

OpenStudy (anonymous):

201

OpenStudy (anonymous):

Ok. x + y = 37 5x + 6y = 201 x represents the number of 5 lettered words and y represents the number of 6 lettered words.